OpenStudy (anonymous):

Mandy is skating at 5 miles per hour. Nikki is skating at 6 miles per hour and started 1 mile behind Mandy. If their rates stay the same, will Mandy catch up with Nikki? Explain.

OpenStudy (anonymous):

make an equation to please

OpenStudy (anonymous):

if nikki maintains a faster rate, given enough time, what do you suppose will happen ?

OpenStudy (anonymous):

she would run faster

OpenStudy (anonymous):

rate = distance/time equivalently distance = rate * time Mandy's distance = 5t Nikki's distance = 6(t - 1) solve for t and that is when the two people will occupy the same distance (i.e., catch up)

OpenStudy (anonymous):

can you help me im confused @binarymimic

OpenStudy (anonymous):

Let Mandy's rate of speed = 5 miles/hour The distance Mandy travels is speed * time = 5 * t = 5t Nikki's rate of speed = 6 miles/hour The distance Nikki travels is speed * time = 6*t = 6t But Nikki started 1 mile behind How long does it take Nikki to travel 1 mile? 6t = 1, t = 1/6 hour So if Nikki starts 1 mile before Mandy, that is the same as saying Nikki starts 1/6 hours before Mandy or Nikki's distance = 6(t - 1/6) = 6t - 1 miles Set up equations 5t = 6t - 1 And this will tell you at what time the two cross paths